# Comprehensive Evaluation of Prototype Neural Networks

### v2

@@ Abstract
Prototype models are an important method for explainable artificial intelligence (XAI) and interpretable machine learning. In this paper, we perform an in-depth analysis of a set of prominent prototype models including ProtoPNet, ProtoPool and PIPNet. For their assessment, we apply a comprehensive set of metrics. In addition to applying standard metrics from literature, we propose several new metrics to further complement the analysis of model interpretability. In our experimentation, we apply the set of prototype models on a diverse set of datasets including fine-grained classification, Non-IID settings and multi-label classification to further contrast the performance. Furthermore, we also provide our code as an open-source library (https://github.com/uos-sis/quanproto), which facilitates simple application of the metrics itself, as well as extensibility – providing the option for easily adding new metrics and models.
keywords:
Interpretable AI, Prototype-based Models, Explainable AI
@@ 1 Introduction
While there have been enormous advances in deep learning [1, 2], the challenge of developing truly interpretable models with accuracy comparable to their black-box counterparts remains unresolved [3, 4, 5]. This is a prominent and active area of research connecting both interpretability and explainability of the respective models [6]. In particular, in the domain of image-related tasks like classification, non-interpretable models typically outperform most other models or techniques. However, in high-risk domains often encountered in industry or the medical sectors, explainability and interpretability are key requirements [7, 8], also enabling computational sensemaking [9].
@@
Different post-hoc techniques have been developed in the past, such as SHAP and LIME [10, 11]. These explain the decision of a model by calculating the importance of individual input features and can be applied to a large family of models. In the domain of image classification, saliency maps are a particular prominent subgroup of these model-agnostic methods, e. g., Integrated Gradients, Grad-CAM and other variants explain the attribution of individual or groups of pixels for a given prediction [12, 13]. Although, such model-agnostic techniques are widely used and can provide considerable insight into trained models, they are in themselves only post-hoc methods with several drawbacks, e. g., [14] compared to a completely intrinsically interpretable model.
@@
One possible solution to overcome some of these problems, has been the development of deep part-based prototype models, firstly presented in [15] and [16]. A key advantage is their intrinsic interpretability. Part-based prototype models focus on learning a meaningful embedding space using an often pre-trained black-box network as feature extractor. A key distinction from standard black-box models is the aim to extract features that concentrate on specific sections of the image. This division of the image in the embedding space allows the network to learn prototypes that represent only particular parts of the object. The classification layer then utilizes the presence of a prototype in the image to calculate the prediction. To provide an explanation for a prediction, the prototypes that have the highest evidence are projected back to the input space through a saliency method.
@@
However, users have different experiences with model explanations. Confirmation bias and subjective understanding make it hard to guess the prediction correctly [17] [18]. This shows that these models are not yet reliable enough to use for AI-assisted decision-making. This encouraged us to survey existing assessments of explainability using a well-founded categorization.
@@
Our research is based on and extends the work of Nauta et al. [19]. With an implementation of a substantial portion of the evaluation framework presented in [19], we facilitate a thorough assessment of part-based prototype models. Our contributions are summarized as follows:
@@
    We expand the set of metrics referenced in [19] by proposing several novel metrics designed to enhance the interpretability assessment of part-based prototype models.
@@
    Furthermore, we conducted a comprehensive analysis of three prominent part-based prototype models, i. e., ProtoPNet [16], ProtoPool [20] and PIPNet [21]. We provide new insights and demonstrate the utility and applicability of our library.
@@
    Our evaluation offers a broad perspective on potential application scenarios by utilizing established datasets such as CUB200 [22] and Cars196 [23] while also extending to a more practice-oriented Non-IID context with the NICO [24] dataset and a multi-label classification task focused on the challenge of learning class-independent features from animals using the AWA2[25] dataset.
@@ 2 Related Work
Central to concept-based explanation is the notion of conceptualization [26] and concepts which has already been investigated in different fields, most notably in case-based reasoning [27, 28]. The basic idea is to focus on human understandable higher-order concepts not only related to single features. Here, a recent and closely related direction is concept-based artificial intelligence [29]. Examples of this approach are e. g., concept bottleneck models [30] or concept activation vectors [31]. Another approach, which shares similarities with prototype-based models, is TesNet by Wang et al. [32]. Here, an interpretable embedding space constructed by basis-concepts is learned.
@@
The prototypical part network (ProtoPNet) from Chen et al. [16] is one of the pioneering networks that popularized the concept of part prototype networks in image classification. In addition to the models examined in this article, other networks have been developed that combine the prototype approach with many concepts from the ML field. Nauta et al. [33] varied the classification process, by replacing the fully connected layer at the end of the model through a decision tree (ProtoTree[33]. ProtoPShare [34] uses an iterative post-processing technique to merge prototypes from a ProtoPNet model, effectively creating shared prototypes between classes and reducing the amount of needed prototypes. Xue et al. [35] adapted the idea of ProtoPNet to be used together with vision transformers as feature extractors. Prototype-based deep learning models have also been used for time series [36], graph data [37], or image segmentation tasks [38]. There is also work, which combines deep prototype models with an autoencoder architecture [39]. Other variants of the ProtoPNet architecture, include [40][41][42][43]. Li et al. [44] provide a comprehensive overview of the design choices and prototype formulations that are used in the literature analysing prototype quality relations.
@@
Part-based prototype models have been used in several industrial application settings e. g., to inspect power grids visually [45], in the classification of MRI scans to detect Alzheimer’s disease [46] or X-Ray images to classify Chest X-rays [47]. Carlone et al. [48] propose a prototype-based model to identify breast cancer, highlighting the acceptance of visual explanations.
@@
Besides the success of ProtoPNet and the improvement through architectural changes, there are also limitations and some critique regarding prototypical networks. Elhadri et al. [49] survey general quality issues regarding the interpretability of learned prototypes in these models. Xu-Darme et al. [50] criticize the way ProtoPNet and ProtoTree visualize the learned prototypes and propose to use different saliency methods to overcome this problem. Hoffmann et al. [51] point out that ProtoPNets struggle with adversarial examples or compression artifacts in the input. Nevertheless, Hoffmann et al. do not want to dispute part prototypical networks, but raise the awareness of these problems. Saralajew et al. [52] contribute to this discussion by introducing a method for fast and provable adversarial robustness certification for Nearest Prototype Classifiers (NPCs).
@@
In the literature, there have been several approaches aiming at evaluating different prototypical-part models. These include different evaluation methods as well as specific metrics. The work of Huang et al. [53], e. g., proposes two metrics called consistency score and stability score that focus on the semantic properties of prototypes alongside a new network architecture to improve these metrics. Recently, there has also been work to create frameworks to simplify the utilization of different types of part-based prototype models, cf. ProtoPNeXt [54] and the work on the library CaBRNet [55]. Ma et al. [56], try to improve the visualization of the prototypes using multiple examples from the training set to describe the concept behind a learned prototype. The framework HIVE, created by Kim et al. [47], is designed for the evaluation of explanation techniques. They highlight key issues that are crucial when humans must interpret these explanations. Nauta et al. [57] examine, why the model considers a prototype and a patch of the input image as similar. A comprehensive survey focusing on the interpretability assessment of part-based prototype models is also given by Nauta et al. [19].
@@ 3 Datasets
The first fine-grained dataset is Caltech-UCSD Birds-200-2011 (CUB200) [22], that features a challenging collection of bird species. This dataset is frequently used as a benchmark for part-based prototype methods. The rich annotations allow for a detailed evaluation of the importance of individual object parts in the classification process, enabling a comprehensive assessment of the prototypes. The dataset contains 11,788 images of 200 bird species with segmentation masks, bounding boxes, part locations, and attribute labels.
@@
For our second fine-grained dataset, we selected the Stanford Cars dataset (Cars196), introduced by Krause et al. [23]. Unlike the natural objects in the CUB200 dataset, this dataset focuses on fine-grained classification of manufactured objects, and is similarly used to benchmark part-based prototype methods. It consists of 16,185 images of 196 car models, covering a wide range of vehicle types including sedans, SUVs, coupes, convertibles, pickups, and hatchbacks.
@@
For general object detection in a Non-IID setting, we used the NICO Dataset from He et al. [24]. This dataset includes 10 animal classes and 9 vehicle classes, though our focus is on the animal classes. Each class is further divided into 10 context categories representing various environments, attributes, and activities. The dataset contains 12,980 images of animals. Essentially, this Non-IID setting is a more realistic scenario, which is also new for part-based prototypical models. Hence, we want to apply this setting mainly through different background environments in the training and test sets to assess if the networks naturally promote object-focused prototypes. Since this cannot be achieved using the provided context categories (as few represent distinct environments), we use a simple image analysis inspired by Cheng et al. [58], based on the HSV colour space [59], to split the dataset into training/validation and test sets. This approach effectively creates a Non-IID scenario with varying distributions across training and test data. Further details are given in the implementation.
@@
The last dataset is Animals with Attributes 2 (AWA2) introduced by Xian et al. [25]. This dataset is a benchmark for attribute-based classification, a related model domain to part-based prototype models. The dataset comprises 37,322 images across 50 animal classes, with binary and continuous class attributes. For evaluating the performance of part-based prototype methods in a multi-label classification setting, we focus on the given attributes. The AWA2 dataset includes 85 attributes that are used to describe a class. Since not all attributes describe visible features, we only use a subset, i. e., those corresponding to the 49 visible attributes.
@@ 4 Prototype Networks
Figure 1: A visual representation of a prototype-based neural network. Starting from (A) a deep neural network, such as ResNet, extracting a feature map (B), the model identifies the presence of prototypes (C) by comparing them with feature vectors; generating similarity maps (D). Classification (F) is performed according to the highest similarity scores (E) derived from these maps. Prototypes are visualized (G) using the Prototypical Relevance Propagation (PRP) method. Note cards list the metrics used in our evaluation for each component.
Because this work focuses on a comprehensive set of metrics, we kept our set of models small and selected three prominent networks based on the following criteria: class-specific vs. shared prototypes and explicit vs. indirect prototype representations. We chose ProtoPNet [16], the first popular prototype model, that uses class-specific explicit prototype vectors. We further include the ProtoPool [20] model as a representative for shared and explicit prototype learning. Finally, we chose PIPNet [21] as a recently introduced architecture for the shared and indirect case. To the best of our knowledge, there is currently no model for the class-specific indirect case. Other models with different classification approaches, such as hierarchical classifications or other backbones such as transformers, are further important studies that should be investigated in connection with the interpretability of prototypes. However, due to a lack of implementations of visual methods and complex adjustments for the multi-label scenario, we decided to start with models that can be easily adapted to our training and evaluation setups.
@@
ProtoPNet uses a multistep training process, separating the training of the last layer from the rest of the network. At the beginning, the last layer is initialized using $1$ for weights connecting a prototype to its corresponding class and 0.5-0.5 otherwise. This ensures the learning of class specific prototypes. The first training step optimizes the backbone and prototype layer to learn a meaningful latent space, clustering prototypes from the same class and separating them from prototypes of other classes. This is done using a linear combination of a classification loss and two novel losses, the cluster loss and separation loss $\lambda_{C}\mathcal{L}_{C}+\lambda_{cl}\mathcal{L}_{cl}+\lambda_{sep}\mathcal{L}_{sep}$. In the second step, each prototype is first projected onto the nearest feature vector from its assigned class. The last layer is then fine-tuned to improve classification performance, using only the classification loss and an L1 regularization term to keep the positive reasoning characteristic. Chen et al. [16] observed that some prototypes focus on the background. These prototypes contradict the goal of the model to learn class-specific object parts. To address this, the authors propose a pruning step after training, which removes such prototypes. We use both the original and pruned version in the evaluation to analyse the effectiveness of this approach.
@@
The ProtoPool [20] model by Rymarczyk et al., addresses several limitations of ProtoPNet. It introduces a new similarity function and an automatic, fully differentiable assignment of prototypes to classes, that enables the network to learn shared prototypes. The model itself follows the architecture of ProtoPNet consisting of a CNN backbone with additional convolutional layer to adjust the output shape, a prototype layer, and a fully connected layer. The prototype layer contains a set of $n$ trainable prototypes $\mathbf{P}=\left\{\mathbf{p}\right\}^{n}$ with $\mathbf{p}\in\mathcal{R}^{1\times 1\times D}$, similar to ProtoPNet. A novel part is the assignment matrix with $L$ prototype slots for each class, where each slot is implemented as a distribution $q_{l}\in\mathbb{R}^{N}$ over all available prototypes. The similarity maps $\mathbf{A}=\left\{\mathbf{a}\right\}^{n}$ are computed using the same equation 1 as in ProtoPNet. To enhance prototype locality, the authors introduce a focal similarity function, which computes the difference between the maximum similarity and the average similarity.
@@
Nauta et al. introduced PIPNet [21], which incorporates two novel loss functions based on CARL (Consistent Assignment for Representation Learning) [60] to learn prototypes that more accurately align with real object parts. Unlike ProtoPNet or ProtoPool, PIPNet does not use explicit prototype vectors. This removes the need to project prototypes onto nearby latent patches. The architecture includes a CNN backbone (without additional convolutional layers), followed by a softmax layer and a fully connected layer. The feature map $\mathbf{z}\in\mathcal{R}^{H\times W\times D}$ from the backbone is treated as $D$ two-dimensional $(H\times W)$ prototype similarity maps. The softmax operation across the $D$ dimension encourages a patch from the feature map $\boldsymbol{\tilde{z}}_{h,w}$, to belong to exactly one prototype similarity map. The similarity scores are again computed via a max-pooling operation and used in the fully connected layer for classification.

### v3

@@ Abstract
Prototype models are an important method for explainable artificial intelligence (XAI) and interpretable machine learning. In this paper, we perform an in-depth analysis of a set of prominent prototype models including ProtoPNet, ProtoPool and PIPNet. For their assessment, we apply a comprehensive set of metrics. In addition to applying standard metrics from literature, we propose several new metrics to further complement the analysis of model interpretability. In our experimentation, we apply the set of prototype models on a diverse set of datasets including fine-grained classification, Non-IID settings and multi-label classification to further contrast the performance. Furthermore, we also provide our source code as an open-source library (https://github.com/uos-sis/quanproto), which facilitates simple application of the metrics itself, as well as extensibility – providing the option for easily adding new metrics and models.
@@ 1 Introduction
While there have been enormous advances in deep learning Wang et al. [2020], Pouyanfar et al. [2018], the challenge of developing truly interpretable models with accuracy comparable to their black-box counterparts remains unresolved Rudin et al. [2022], Du et al. [2019], Zhang et al. [2021]. This is a prominent and active area of research connecting both interpretability and explainability of the respective models Atzmueller et al. [2024]. In particular, in the domain of image-related tasks like classification, non-interpretable models typically outperform most other models or techniques. However, in high-risk domains often encountered in industry or the medical sectors, explainability and interpretability are key requirements Rudin [2019], Stiglic et al. [2020], also enabling computational sensemaking Atzmueller [2017].
@@
Different post-hoc techniques have been developed in the past, such as SHAP and LIME Lundberg and Lee [2017], Ribeiro et al. [2016]. These explain the decision of a model by calculating the importance of individual input features and can be applied to a large family of models. In the domain of image classification, saliency maps are a particular prominent subgroup of these model-agnostic methods, e. g., Integrated Gradients, Grad-CAM and other variants explain the attribution of individual or groups of pixels for a given prediction Qi et al. [2019], Selvaraju et al. [2020]. Although, such model-agnostic techniques are widely used and can provide considerable insight into trained models, they are in themselves only post-hoc methods with several drawbacks, e. g., Adebayo et al. [2018] compared to a completely intrinsically interpretable model.
@@
One possible solution to overcome some of these problems, has been the development of deep part-based prototype models, firstly presented in Li et al. [2018] and Chen et al. [2019]. A key advantage is their intrinsic interpretability. Part-based prototype models focus on learning a meaningful embedding space using an often pre-trained black-box network as feature extractor. A key distinction from standard black-box models is the aim to extract features that concentrate on specific sections of the image. This division of the image in the embedding space allows the network to learn prototypes that represent only particular parts of the object. The classification layer then utilizes the presence of a prototype in the image to calculate the prediction. To provide an explanation for a prediction, the prototypes that have the highest evidence are projected back to the input space through a saliency method.
@@
However, users have different experiences with model explanations. Confirmation bias and subjective understanding make it hard to guess the prediction correctly Kim et al. [2022] Nguyen et al. [2021]. This shows that these models are not yet reliable enough to use for AI-assisted decision-making. This encouraged us to survey existing assessments of explainability using a well-founded categorization.
@@
Our research is based on and extends the work of Nauta et al. Nauta and Seifert [2023]. With an implementation of a substantial portion of the evaluation framework presented in Nauta and Seifert [2023], we facilitate a thorough assessment of part-based prototype models. Our contributions are summarized as follows:
@@
    We expand the set of metrics referenced in Nauta and Seifert [2023] by proposing several novel metrics designed to enhance the interpretability assessment of part-based prototype models.
@@
    Furthermore, we conducted a comprehensive analysis of three prominent part-based prototype models, i. e., ProtoPNet Chen et al. [2019], ProtoPool Rymarczyk et al. [2022] and PIPNet Nauta et al. [2023a]. We provide new insights and demonstrate the utility and applicability of our library.
@@
    Our evaluation offers a broad perspective on potential application scenarios by utilizing established datasets such as CUB200 Wah et al. [2011] and Cars196 Krause et al. [2013] while also extending to a more practice-oriented Non-IID context with the NICO He et al. [2021] dataset and a multi-label classification task focused on the challenge of learning class-independent features from animals using the AWA2Xian et al. [2018] dataset.
@@ 2 Related Work
Central to concept-based explanation is the notion of conceptualization Schank [1975] and concepts which has already been investigated in different fields, most notably in case-based reasoning Kolodner [1992], Sørmo et al. [2005]. The basic idea is to focus on human understandable higher-order concepts not only related to single features. Here, a recent and closely related direction is concept-based artificial intelligence Poeta et al. [2023]. Examples of this approach are e. g., concept bottleneck models Koh et al. [2020] or concept activation vectors Kim et al. [2018]. Another approach, which shares similarities with prototype-based models, is TesNet by Wang et al. Wang et al. [2021]. Here, an interpretable embedding space constructed by basis-concepts is learned.
@@
The prototypical part network (ProtoPNet) from Chen et al. Chen et al. [2019] is one of the pioneering networks that popularized the concept of part prototype networks in image classification. In addition to the models examined in this article, other networks have been developed that combine the prototype approach with many concepts from the ML field. Nauta et al. Nauta et al. [2021a] varied the classification process, by replacing the fully connected layer at the end of the model through a decision tree (ProtoTreeNauta et al. [2021a]. ProtoPShare Rymarczyk et al. [2021] uses an iterative post-processing technique to merge prototypes from a ProtoPNet model, effectively creating shared prototypes between classes and reducing the amount of needed prototypes. Xue et al. Xue et al. [2022] adapted the idea of ProtoPNet to be used together with vision transformers as feature extractors. Prototype-based deep learning models have also been used for time series Ming et al. [2019], graph data Zhang et al. [2022], or image segmentation tasks Sacha et al. [2023a]. There is also work, which combines deep prototype models with an autoencoder architecture Gautam et al. [2022]. Other variants of the ProtoPNet architecture, include Pach et al. [2024]van der Klis et al. [2023]Carmichael et al. [2024]Donnelly et al. [2022]. Li et al. Li et al. [2024] provide a comprehensive overview of the design choices and prototype formulations that are used in the literature analysing prototype quality relations.
@@
Part-based prototype models have been used in several industrial application settings e. g., to inspect power grids visually Stefenon et al. [2022], in the classification of MRI scans to detect Alzheimer’s disease Mohammadjafari et al. [2021] or X-Ray images to classify Chest X-rays Kim et al. [2021]. Carlone et al. Carloni et al. [2022] propose a prototype-based model to identify breast cancer, highlighting the acceptance of visual explanations.
@@
Besides the success of ProtoPNet and the improvement through architectural changes, there are also limitations and some critique regarding prototypical networks. Elhadri et al. Elhadri et al. [2025] survey general quality issues regarding the interpretability of learned prototypes in these models. Xu-Darme et al. Xu-Darme et al. [2023a] criticize the way ProtoPNet and ProtoTree visualize the learned prototypes and propose to use different saliency methods to overcome this problem. Hoffmann et al. Hoffmann et al. [2021] point out that ProtoPNets struggle with adversarial examples or compression artifacts in the input. Nevertheless, Hoffmann et al. do not want to dispute part prototypical networks, but raise the awareness of these problems. Saralajew et al. Saralajew et al. [2020] contribute to this discussion by introducing a method for fast and provable adversarial robustness certification for Nearest Prototype Classifiers (NPCs).
@@
In the literature, there have been several approaches aiming at evaluating different prototypical-part models. These include different evaluation methods as well as specific metrics. The work of Huang et al. Huang et al. [2023], e. g., proposes two metrics called consistency score and stability score that focus on the semantic properties of prototypes alongside a new network architecture to improve these metrics. Recently, there has also been work to create frameworks to simplify the utilization of different types of part-based prototype models, cf. ProtoPNeXt Willard et al. [2024] and the work on the library CaBRNet Xu-Darme et al. [2024]. Ma et al. Ma et al. [2024], try to improve the visualization of the prototypes using multiple examples from the training set to describe the concept behind a learned prototype. The framework HIVE, created by Kim et al. Kim et al. [2021], is designed for the evaluation of explanation techniques. They highlight key issues that are crucial when humans must interpret these explanations. Nauta et al. Nauta et al. [2021b] examine, why the model considers a prototype and a patch of the input image as similar. A comprehensive survey focusing on the interpretability assessment of part-based prototype models is also given by Nauta et al. Nauta and Seifert [2023].
@@ 3 Datasets
The first fine-grained dataset is Caltech-UCSD Birds-200-2011 (CUB200) Wah et al. [2011], that features a challenging collection of bird species. This dataset is frequently used as a benchmark for part-based prototype methods. The rich annotations allow for a detailed evaluation of the importance of individual object parts in the classification process, enabling a comprehensive assessment of the prototypes. The dataset contains 11,788 images of 200 bird species with segmentation masks, bounding boxes, part locations, and attribute labels.
@@
For our second fine-grained dataset, we selected the Stanford Cars dataset (Cars196), introduced by Krause et al. Krause et al. [2013]. Unlike the natural objects in the CUB200 dataset, this dataset focuses on fine-grained classification of manufactured objects, and is similarly used to benchmark part-based prototype methods. It consists of 16,185 images of 196 car models, covering a wide range of vehicle types including sedans, SUVs, coupes, convertibles, pickups, and hatchbacks.
@@
For general object detection in a Non-IID setting, we used the NICO Dataset from He et al. He et al. [2021]. This dataset includes 10 animal classes and 9 vehicle classes, though our focus is on the animal classes. Each class is further divided into 10 context categories representing various environments, attributes, and activities. The dataset contains 12,980 images of animals. Essentially, this Non-IID setting is a more realistic scenario, which is also new for part-based prototypical models. Hence, we want to apply this setting mainly through different background environments in the training and test sets to assess if the networks naturally promote object-focused prototypes. Since this cannot be achieved using the provided context categories (as few represent distinct environments), we use a simple image analysis inspired by Cheng et al. Cheng et al. [2001], based on the HSV colour space contributors [2024], to split the dataset into training/validation and test sets. This approach effectively creates a Non-IID scenario with varying distributions across training and test data. Further details are given in the implementation.
@@
The last dataset is Animals with Attributes 2 (AWA2) introduced by Xian et al. Xian et al. [2018]. This dataset is a benchmark for attribute-based classification, a related model domain to part-based prototype models. The dataset comprises 37,322 images across 50 animal classes, with binary and continuous class attributes. For evaluating the performance of part-based prototype methods in a multi-label classification setting, we focus on the given attributes. The AWA2 dataset includes 85 attributes that are used to describe a class. Since not all attributes describe visible features, we only use a subset, i. e., those corresponding to the 49 visible attributes.
@@ 4 Prototype Networks
Figure 1: A visual representation of a prototype-based neural network. Starting from (A) a deep neural network, such as ResNet, extracting a feature map (B), the model identifies the presence of prototypes (C) by comparing them with feature vectors; generating similarity maps (D). Classification (F) is performed according to the highest similarity scores (E) derived from these maps. Prototypes are visualized (G) using the Prototypical Relevance Propagation (PRP) method. Note cards list the metrics that focus on that specific component. Some metrics are used in multiple evaluation techniques, which is later indicated by a subscript, see table 1
Because this work focuses on a comprehensive set of metrics, we kept our set of models small and selected three prominent networks based on the following criteria: class-specific vs. shared prototypes and explicit vs. indirect prototype representations. We chose ProtoPNet Chen et al. [2019], the first popular prototype model, that uses class-specific explicit prototype vectors. We further include the ProtoPool Rymarczyk et al. [2022] model as a representative for shared and explicit prototype learning. Finally, we chose PIPNet Nauta et al. [2023a] as a recently introduced architecture for the shared and indirect case. To the best of our knowledge, there is currently no model for the class-specific indirect case. Other models with different classification approaches, such as hierarchical classifications or other backbones such as transformers, are further important studies that should be investigated in connection with the interpretability of prototypes. However, due to a lack of implementations of visual methods and complex adjustments for the multi-label scenario, we decided to start with models that can be easily adapted to our training and evaluation setups.
@@
ProtoPNet uses a multistep training process, separating the training of the last layer from the rest of the network. At the beginning, the last layer is initialized using $1$ for weights connecting a prototype to its corresponding class and 0.5-0.5 otherwise. This ensures the learning of class specific prototypes. The first training step optimizes the backbone and prototype layer to learn a meaningful latent space, clustering prototypes from the same class and separating them from prototypes of other classes. This is done using a linear combination of a classification loss and two novel losses, the cluster loss and separation loss $\lambda_{C}\mathcal{L}_{C}+\lambda_{cl}\mathcal{L}_{cl}+\lambda_{sep}\mathcal{L}_{sep}$. In the second step, each prototype is first projected onto the nearest feature vector from its assigned class. The last layer is then fine-tuned to improve classification performance, using only the classification loss and an L1 regularization term to keep the positive reasoning characteristic. Chen et al. Chen et al. [2019] observed that some prototypes focus on the background. These prototypes contradict the goal of the model to learn class-specific object parts. To address this, the authors propose a pruning step after training, which removes such prototypes. We use both the original and pruned version in the evaluation to analyse the effectiveness of this approach.
@@
The ProtoPool Rymarczyk et al. [2022] model by Rymarczyk et al., addresses several limitations of ProtoPNet. It introduces a new similarity function and an automatic, fully differentiable assignment of prototypes to classes, that enables the network to learn shared prototypes. The model itself follows the architecture of ProtoPNet consisting of a CNN backbone with additional convolutional layer to adjust the output shape, a prototype layer, and a fully connected layer. The prototype layer contains a set of $n$ trainable prototypes $\mathbf{P}=\left\{\mathbf{p}\right\}^{n}$ with $\mathbf{p}\in\mathcal{R}^{1\times 1\times D}$, similar to ProtoPNet. A novel part is the assignment matrix with $L$ prototype slots for each class, where each slot is implemented as a distribution $q_{l}\in\mathbb{R}^{N}$ over all available prototypes. The similarity maps $\mathbf{A}=\left\{\mathbf{a}\right\}^{n}$ are computed using the same equation 1 as in ProtoPNet. To enhance prototype locality, the authors introduce a focal similarity function, which computes the difference between the maximum similarity and the average similarity.
@@
Nauta et al. introduced PIPNet Nauta et al. [2023a], which incorporates two novel loss functions based on CARL (Consistent Assignment for Representation Learning) Silva and Rivera [2022] to learn prototypes that more accurately align with real object parts. Unlike ProtoPNet or ProtoPool, PIPNet does not use explicit prototype vectors. This removes the need to project prototypes onto nearby latent patches. The architecture includes a CNN backbone (without additional convolutional layers), followed by a softmax layer and a fully connected layer. The feature map $\mathbf{z}\in\mathcal{R}^{H\times W\times D}$ from the backbone is treated as $D$ two-dimensional $(H\times W)$ prototype similarity maps. The softmax operation across the $D$ dimension encourages a patch from the feature map $\boldsymbol{\tilde{z}}_{h,w}$, to belong to exactly one prototype similarity map. The similarity scores are again computed via a max-pooling operation and used in the fully connected layer for classification.